OpenStudy (tomektomeknyc):

If theta = pi/4 , Beta = pi/2 , sec (theta- beta) = a) - csc pi/4 b) sec pi/4 c) csc pi/4 d) - sec pi/4

OpenStudy (john_es):

I would calculate first, the difference, theta-beta=pi/4-pi/2=-pi/4 Then I would use that cos (-a)=cos (a) And sec(a)=1/cos(a)

OpenStudy (john_es):

I think you can follow from this point.

OpenStudy (tomektomeknyc):

I m sorry, I dont understand this part: pi/4 - pi/2 = -pi/4

OpenStudy (john_es):

theta=pi/4 beta=pi/2 theta-beta=pi/4-pi/2=pi*(1/4-1/2)=pi*(1/4-2/4)=pi*(-1/4)=-pi/4 Hope it helps.

OpenStudy (tomektomeknyc):

Thank you

OpenStudy (tomektomeknyc):

I think it will be answer b